OpenStudy (anonymous):

what percent of 496 is 124

OpenStudy (austinl):

\[\frac{496}{124} = x\%\]

OpenStudy (lncognlto):

Um, I think its (124/496)*100 = x%

OpenStudy (lncognlto):

@Luigi0210

OpenStudy (austinl):

Yes, silly error. My apologies.

OpenStudy (austinl):

\[\frac{124}{496}\times 100 = x\%\]

OpenStudy (luigi0210):

Haha, I thought Austin lost his touch for a second there

OpenStudy (luigi0210):

\[\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile } \]

OpenStudy (austinl):

I started out that way, and my brain was saying "nooooo, thats not how you do it you silly potato." Should have went with my gut.

OpenStudy (lncognlto):

LOL austin. And yes, Welcome to OS